OpenStudy (anonymous):

The manager of a stadium wants to know how many baseball season ticket holders also have season tickets for other sports. There are 628 season ticket holders. She randomly selects 75 of those ticket holders and finds that 21 of them have season tickets for other sports. Based on this sample, about how many of the baseball season ticket holders have tickets for other sports? a. 224 b. 176 c. 157 d. 132 my answer: b. 176 b/c i did this: 21/75*628=175.84 which is about 176... am i right??

OpenStudy (campbell_st):

75/628 = 21/x x = (21 * 628)/75

OpenStudy (anonymous):

i still got 175.84... so is my answer right??? the answer is 176 then??

OpenStudy (campbell_st):

looks good

OpenStudy (anonymous):

thanks @campbell_st :)

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

yes it did help, thank you!